# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks

### v2

@@ Abstract
Deep Neural Networks (DNNs) are increasingly deployed across a wide range of applications, from image classification to autonomous driving. However, ensuring their reliability remains a challenge and in many situations, alternative models with similar functionality and accuracy levels are available. Traditional accuracy-based evaluations often fail to capture behavioral differences between such models, particularly when testing datasets are limited, making it challenging to select or optimally combine models.
Differential testing addresses this limitation by generating test inputs that expose discrepancies in the behavior of DNN models. However, existing differential testing approaches face significant limitations: many rely on access to model internals or are constrained by the availability of seed inputs, limiting their generalizability and effectiveness. In response to these challenges, we propose DiffGAN, a black-box test generation approach for differential testing of DNN models. Our method relies on a Generative Adversarial Network (GAN) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to generate diverse and valid triggering inputs that effectively reveal behavioral discrepancies between models. Our method employs two custom fitness functions, one focused on diversity and the other on divergence, to guide the exploration of the GAN input space and identify discrepancies between the models’ outputs. By strategically searching the GAN input space, we show that DiffGAN can effectively generate inputs with specific features that trigger differences in behavior for the models under test.
Unlike traditional white-box methods, DiffGAN does not require access to the internal structure of the models, which makes it applicable to a wider range of situations. We evaluate DiffGAN on a benchmark comprising eight pairs of DNN models trained on two widely used image classification datasets. Our results demonstrate that DiffGAN significantly outperforms a state-of-the-art (SOTA) baseline, generating four times more triggering inputs, with higher diversity and validity, within the same testing budget. Furthermore, we show that the generated input can be used to improve the accuracy of a machine learning-based model selection mechanism, which dynamically selects the best performing model based on input characteristics and can thus be used as a smart model output voting mechanism when using alternative models together.
@@ 1 Introduction
These challenges underscores the need for differential testing, an approach that focuses on generating triggering inputs—test inputs specifically designed to expose behavioral discrepancies between DNN models under test. Differential testing becomes particularly crucial in scenarios such as model selection, model compression, regression analysis, and the use of model ensembles. In the case of model selection, organizations often need to choose from several models that have been trained for the same task, possibly with different architectures, training data, or optimization techniques.
@@
Although these models may achieve similar accuracy on a given test dataset, their performance can vary significantly under different operational conditions. Differential testing helps identify these variations by generating inputs that expose where one model might outperform the others, providing deeper insights that go beyond simple accuracy metrics. Similarly, in model compression, large DNN models are often compressed using techniques such as quantization, pruning, or knowledge distillation to make them suitable for deployment in resource-constrained environments, such as mobile devices or embedded systems. Differential testing is essential here to ensure that the compressed models retain the accuracy of the original uncompressed versions. By generating inputs that highlight behavioral differences between the compressed and original models, differential testing allows developers to better understand the trade-offs involved in compression and to ensure that the model’s critical performance characteristics are preserved. For regression analysis, when DNN models are updated or re-trained, differential testing identifies and generates inputs where the new version performs worse than the original or vice versa, ensuring that updates do not compromise the accuracy of the model. Last, when using ensembles of models, a voting mechanism is required and one needs to be able to learn when to trust which model’s output.
@@
However, the differential testing process presents several challenges. First, the input subspace where highly accurate models mispredict is often extremely limited, within an extremely large input space. This makes it difficult to find the specific inputs that will trigger behavioral differences, especially in models designed to perform well across a wide range of scenarios. Efficient and effective techniques are thus required to reveal meaningful such discrepancies. Second, many models are treated as black-boxes, where only the inputs and outputs are visible without access to the internal workings of the model. This lack of access complicates the generation of test inputs since the internal decision-making processes remain hidden. Third, limited testing budgets constitute another constraint, particularly in scenarios where collecting, labeling, and validating test input is time-consuming and resource-intensive. In cases where models are deployed in environments that require expensive simulators or manual validation, testing costs can quickly escalate, further limiting the resources available for comprehensive testing.
@@
Finally, ensuring the validity of the generated inputs in differential testing poses a significant challenge. Invalid inputs can lead to mispredictions and skew models’ performance assessments. Human feedback is often required to verify the contextual relevance and validity of test inputs, adding complexity and making the process more resource intensive. This manual validation step is difficult to scale, further complicating efforts to evaluate true performance discrepancies between models effectively.
@@
Several test generation approaches for the differential testing of DNN models have been proposed in the literature [4, 5, 6, 7, 8, 9, 10], but they often come with key limitations. Many of these methods are white-box [4, 5, 6, 7], requiring access to the internal structure and parameters of the DNN models under test. This restricts their applicability to proprietary models, where such internal access is typically unavailable due to privacy or intellectual property concerns. Some other approaches [6, 7] assume structural similarities between the models under test, limiting their applicability on DNN models of different architecture. Black-box differential testing methods [8, 9, 10], on the other hand, often rely on customized mutation operators applied to seed inputs to generate triggering inputs. However, these methods are highly dependent on the availability and quality of the seed data, restricting the generation process to variations of existing inputs.
@@
In response to these challenges, we propose DiffGAN, a black-box test generation approach tailored for the differential testing of DNN models with similar accuracy levels. Our approach is specific to DNN models used in image classification, a domain in which such models are widely adopted. DiffGAN leverages Generative Adversarial Networks (GANs) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to generate diverse and valid triggering inputs that effectively reveal behavioral disagreements between the models under test. The use of a GAN model for this purpose stems from its proven ability in generating realistic and novel images that conform to the original input domain. This is achieved by exploring the latent space of possible inputs. To effectively guide the generation process, we rely on NSGA-II to effectively search the GAN’s latent space and generate diverse triggering images. The genetic search employs two fitness functions, focusing on maximizing (1) diversity to avoid the generation of redundant triggering images, and (2) divergence to maximize the behavioral disagreements between the DNN models. Since DiffGAN operates as a black-box method, it does not require access to the internal structure of the models or assume structural similarities between them, making it suitable for testing any pair of DNN models. DiffGAN is focused on image classification problems, for which Deep Learning is widely applied across many applications.
@@
To evaluate DiffGAN, we constructed a benchmark consisting of eight pairs of models (16 DNN models) trained on two widely used image classification datasets. We compared its performance to DRfuzz [10], a SOTA differential testing baseline assuming the same testing budget. Our experimental results demonstrate that DiffGAN significantly outperforms the baseline, generating on average four times more triggering inputs. Moreover, these inputs include a greater number of valid images and exhibit higher diversity, enabling a more comprehensive assessment of the models’ behaviors. We further demonstrate an application of the triggering inputs generated by DiffGAN through the development of a machine learning-based selection mechanism. This system uses a machine learning model to dynamically select between two classifiers, predicting which one will perform better based on the input characteristics. Our results show that training the selection model with inputs generated by DiffGAN leads to significantly more accurate selections compared to those trained with inputs from the baseline approach.
@@ 2 Problem Definition
In this paper, we propose a test generation approach for the differential testing of DNN models to reveal behavioral disagreements between DNN models that exhibit similar accuracy levels.
@@ 2.1 Motivations
Differential testing is crucial across a variety of practical applications such as model selection, voting systems, model compression, model regression, and detecting model reuse [8, 9, 11, 10, 12]. For example, consider the context of model selection. In practice, organizations often face the challenge of choosing the most suitable model during the training phase, prior to inference time. This decision process involves choosing from multiple DNN models, each one possibly trained by different providers for the same task, using different model architectures or training datasets, and demonstrating similar accuracy on a testing dataset. In such scenario, the ability to generate inputs that clearly indicate differences in model behavior is invaluable to guide model selection.
@@
Differential testing also plays a pivotal role in identifying model regression [10], which entails detecting discrepancies between iterations of a deep learning model (original versus updated versions). For instance, an update aimed at enhancing the overall accuracy of a DNN model might inadvertently introduce inaccuracies under specific input conditions. To mitigate this, triggering test inputs are essential, to systematically assess changes in accuracy across updates and thus ensure they do not lead to poorer accuracy in critical conditions.
Additionally, differential testing is useful for detecting model reuse [12]. In industries where proprietary models are crucial assets, it is important to verify that deployed models are original. Differential testing can identify if a model presented as new is actually derived from or modified from an existing model, thus protecting intellectual property and ensuring the authenticity of models.
In conclusion, while differential testing has broad applications, our focus in this work is specifically on model selection. Unlike other contexts such as model compression, regression detection, or model reuse, the model selection scenario presents unique challenges. Here, the goal is not to evaluate how a single model evolves or changes, but rather to compare multiple models developed independently for the same task. These models may exhibit similar accuracy levels but differ in training data, structure, and optimization methods. Focusing on model selection allows us to explore how differential testing can expose behavioral differences between DNN models. This focus is critical because, in practice, selecting the most appropriate model, whether after training or at inference time, can directly impact the systems using these models.
@@ 2.2 Challenges
Generating triggering inputs to reveal behavioral disagreements between DNN models under test introduces three main challenges. First, uncovering discrepancies in highly-accurate models is inherently difficult due to the very small input subspace leading to mispredictions. Because they are designed to perform well under a wide range of conditions, we require effective and efficient techniques to identify behavioral disagreements. Second, in many real-world scenarios, we do not have access to the internal information of DNN models. Indeed, such models are often proprietary and deployed on servers, accessible only through cloud APIs. This limitation constrains us to using black-box techniques to generate triggering inputs as we only have access to the inputs and outputs of the DNN models. Finally, the testing budget is often limited, particularly when models (1) require extensive manual labor to label and validate triggering inputs, or (2) operate on expensive simulators or physical devices. These simulators usually provide a controlled environment that replicates true operational scenarios, allowing for rigorous testing without the costs and risks associated with testing models deployed on expensive physical devices. Within a realistic testing budget, we must effectively generate triggering inputs that reveal diverse behavioral disagreements between the models under test.
@@ 2.3 Assumptions
No Access to the Training Data: We assume that we do not have access to the original training datasets of the models under test. This is a common scenario in many real-world situations where models are proprietary or when model testing is provided as a service for example. Consequently, our ability to generate inputs and interpret model behavior is based solely on observations from models’ outputs and the currently available testing data, rather than insights from the training process.
Similar Accuracy Levels: We assume that each pair of models under test yields similar and high accuracy on the currently available testing dataset. Such generation warrants the effective generation of triggering inputs to reveal behavioral disagreement between the DNN models, to better assess their respective capabilities and thus enable their use in the various practical applications mentioned above.
@@ 3.3.1 Individuals and Initial Population
In genetic search, the initial population contains individuals consisting of a set of elements called genes. These genes are connected and form an individual also referred to as a solution. In our context, the initial population contains 100 randomly generated latent vectors, each of size 100, as recommended in the literature [36]. Each dimension of a latent vector in the initial population is independently generated from a uniform distribution over the range $[-1,1]$, thereby ensuring a diverse and unbiased initialization in the search space.
@@ 3.3.2 Fitness Functions
This fitness function estimates the divergence between the output probability vectors of the models under test, when fed with the same input image generated by the trained GAN. More specifically, the fitness function takes as input a GAN latent vector $Z$ responsible for generating the candidate test input image. We feed the models under test with such image and compute the Euclidean distance between the resulting output probability vectors $Y_{m1}$ and $Y_{m2}$. The goal of maximizing this distance is to increase the probability of selecting latent vectors that lead to the generation of input images revealing significant behavioral disagreements between the models under test. The Euclidean distance serves as a straightforward measure of the magnitude of discrepancy between model predictions, capturing how far apart the models’ outputs are in the probability space. The choice of Euclidean distance over other metrics, such as KL divergence [37], is motivated by its efficiency in terms of computation cost.
@@
$\begin{split}\max\limits_{z\in\mathbb{P}}\>\>&Fitness(z)=(Divergence(z),Diversity(z))\\ \end{split}$ (3)
@@ 3.4.2 Mutation
Our genetic search also employs a mutation operator to further explore the search space. Following existing works that use genetic algorithms in GAN latent space [28, 40], we also adopt polynomial mutation [42] to address the continuous nature of the latent space.
@@ 3.5 Search Algorithm
Input: Initial population of individuals ( $ind$in $P$ ), maximum running time ($T_{max}$), trained GAN, two DNNs under test
@@ 14             $\text{image}_{\text{gen}}\xleftarrow{}\text{GAN}(ind)$;
15            $[OP1,OP2]\xleftarrow{}\text{Model1}(\text{image}_{\text{gen}}),\text{Model2}(\text{image}_{\text{gen}})$
@@ 21       $gen=gen+1$;
We set the NSGA-II parameters in our search algorithm as follows. The crossover rate $CrossRate$ is equal to 90% and the mutation rate $MutationRate$ is equal to 10%, given recommended values in the literature [44, 45, 46]. Finally, we used Google Colab and the Pymoo library [47] to implement our genetic search.

### v3

@@ Abstract
Deep Neural Networks (DNNs) are increasingly deployed across a wide range of applications, from image classification to autonomous driving. However, ensuring their reliability remains a challenge, and in many situations, alternative models with similar functionality and accuracy levels are available. Traditional accuracy-based evaluations often fail to capture behavioral differences between such models, particularly when testing datasets are limited, making it challenging to select or optimally combine models.
Differential testing addresses this limitation by generating test inputs that expose discrepancies in the behavior of DNN models. However, existing differential testing approaches face significant limitations: many rely on access to model internals or are constrained by the availability of seed inputs, limiting their generalizability and effectiveness. In response to these challenges, we propose DiffGAN, a black-box test generation approach for differential testing of DNN models. Our approach, though adaptable to other domains, is specific to DNN models for image classification tasks, a highly prevalent application area. Our method relies on a Generative Adversarial Network (GAN) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to generate diverse and valid triggering inputs that effectively reveal behavioral discrepancies between models. Our method employs two custom fitness functions, one focused on diversity and the other on divergence, to guide the exploration of the GAN input space and identify discrepancies between the models’ outputs. By strategically searching the GAN input space, we show that DiffGAN can effectively generate inputs with specific features that trigger differences in behavior for the models under test. Unlike traditional white-box methods, DiffGAN does not require access to the internal structure of the models, which makes it applicable to a wider range of situations. We evaluate DiffGAN on a benchmark comprising eight pairs of DNN models trained on two widely used image classification datasets. Our results demonstrate that DiffGAN significantly outperforms a state-of-the-art (SOTA) baseline, generating four times more triggering inputs, with higher diversity and validity, within the same testing budget. Furthermore, we show that the generated input can be used to improve the accuracy of a machine learning-based model selection mechanism, which dynamically selects the best-performing model based on input characteristics and can thus be used as a smart model output voting mechanism when using alternative models together.
@@ 1 Introduction
These challenges underscore the need for differential testing, an approach that focuses on generating triggering inputs—test inputs specifically designed to expose behavioral discrepancies between DNN models under test. Differential testing becomes particularly crucial in scenarios such as model selection, model compression, regression analysis, and the use of model ensembles. In the case of model selection, organizations often need to choose from several models that have been trained for the same task, possibly with different architectures, training data, or optimization techniques.
@@
Although these models may achieve similar accuracy on a given test dataset, their performance can vary significantly under different operational conditions. Differential testing helps identify these variations by generating inputs that expose where one model might outperform the others, providing deeper insights that go beyond simple accuracy metrics. Similarly, in model compression, large DNN models are often compressed using techniques such as quantization, pruning, or knowledge distillation to make them suitable for deployment in resource-constrained environments, such as mobile devices or embedded systems. Differential testing is essential here to ensure that the compressed models retain the accuracy of the original uncompressed versions. By generating inputs that highlight behavioral differences between the compressed and original models, differential testing allows developers to understand better the trade-offs involved in compression and to ensure that the model’s critical performance characteristics are preserved. For regression analysis, when DNN models are updated or re-trained, differential testing identifies and generates inputs where the new version performs worse than the original or vice versa, ensuring that updates do not compromise the accuracy of the model. Last, when using ensembles of models, a voting mechanism is required, and one needs to be able to learn when to trust which model’s output.
@@
However, the differential testing process presents several challenges. First, the input subspace where highly accurate models mispredict is often extremely limited, within a huge input space. This makes it challenging to find the specific inputs that will trigger behavioral differences, especially in models designed to perform well across a wide range of scenarios. Efficient and effective techniques are thus required to reveal meaningful such discrepancies. Second, many models are treated as black-boxes, where only the inputs and outputs are visible without access to the internal workings of the model. This lack of access complicates the generation of test inputs since the internal decision-making processes remain hidden. Third, limited testing budgets constitute another constraint, particularly in scenarios where collecting, labeling, and validating test input is time-consuming and resource-intensive. In cases where models are deployed in environments that require expensive simulators or manual validation, testing costs can quickly escalate, further limiting the resources available for comprehensive testing.
@@
Finally, ensuring the validity of the generated inputs in differential testing poses a significant challenge. Invalid inputs can lead to mispredictions and skew models’ performance assessments. Human feedback is often required to verify the contextual relevance and validity of test inputs, adding complexity and making the process more resource-intensive. This manual validation step is difficult to scale, further complicating efforts to evaluate true performance discrepancies between models effectively.
@@
Several test generation approaches for the differential testing of DNN models have been proposed in the literature [4, 5, 6, 7, 8, 9, 10], but they often come with key limitations. Many of these methods are white-box [4, 5, 6, 7], requiring access to the internal structure and parameters of the DNN models under test. This restricts their applicability to proprietary models, where such internal access is typically unavailable due to privacy or intellectual property concerns. Some other approaches [6, 7] assume structural similarities between the models under test, limiting their applicability on DNN models of different architectures. Black-box differential testing methods [8, 9, 10], on the other hand, often rely on customized mutation operators applied to seed inputs to generate triggering inputs. However, these methods are highly dependent on the availability and quality of the seed data, restricting the generation process to variations of existing inputs.
@@
In response to these challenges, we propose DiffGAN, a black-box test generation approach tailored for the differential testing of DNN models with similar accuracy levels. Our approach is specific to DNN models used in image classification, a domain in which such models are widely applied. DiffGAN leverages Generative Adversarial Networks (GANs) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to generate diverse and valid triggering inputs that effectively reveal behavioral disagreements between the models under test. The use of a GAN model for this purpose stems from its proven ability to generate realistic and novel images that conform to the original input domain. This is achieved by exploring the latent space of possible inputs. To effectively guide the generation process, we rely on NSGA-II to effectively search the GAN’s latent space and generate diverse triggering images. The genetic search employs two fitness functions, focusing on maximizing (1) diversity to avoid the generation of redundant triggering images, and (2) divergence to maximize the behavioral disagreements between the DNN models. Since DiffGAN operates as a black-box method, it does not require access to the internal structure of the models or assume structural similarities between them, making it suitable for testing any pair of DNN models. DiffGAN is focused on image classification problems, for which Deep Learning is widely applied across many applications.
@@
To evaluate DiffGAN, we constructed a benchmark consisting of eight pairs of models (16 DNN models) trained on two widely used image classification datasets. We compared its performance to DRfuzz [10], a SOTA differential testing baseline assuming the same testing budget. Our experimental results demonstrate that DiffGAN significantly outperforms the baseline, generating on average four times more triggering inputs. Moreover, these inputs include a greater number of valid images and exhibit higher diversity, enabling a more comprehensive assessment of the models’ behaviors.
We further demonstrate an application of the triggering inputs generated by DiffGAN through the development of a machine learning-based selection mechanism. This system uses a machine learning model to dynamically select between two classifiers, predicting which one will perform better based on the input characteristics. Our results show that training the selection model with inputs generated by DiffGAN leads to significantly more accurate selections compared to those trained with inputs from the baseline approach.
@@ 2 Problem Definition
In this paper, we propose a test generation approach for the differential testing of DNN models. The goal is to generate triggering inputs—inputs that expose behavioral discrepancies between DNN models with similar accuracy levels. Though it is adaptable to other domains, our approach focuses on image classification tasks, a prominent application domain for DNN models. Throughout the paper, we define a triggering input as an image that causes the DNN models under test to produce different outputs.
@@ 2.1 Motivations
Differential testing is crucial across a variety of practical applications such as model selection, voting systems, model compression, model regression, and detecting model reuse [8, 9, 11, 10, 12]. For example, consider the context of model selection. In practice, organizations often face the challenge of choosing the most suitable model during the training phase, prior to inference time. This decision process involves choosing from multiple DNN models, each one possibly trained by different providers for the same task, using different model architectures or training datasets, and demonstrating similar accuracy on a testing dataset. In such a scenario, the ability to generate inputs that clearly indicate differences in model behavior is invaluable to guide model selection.
@@
Differential testing also plays a pivotal role in identifying model regression [10], which entails detecting discrepancies between iterations of a deep learning model (original versus updated versions). For instance, an update aimed at enhancing the overall accuracy of a DNN model might inadvertently introduce inaccuracies under specific input conditions. To mitigate this, triggering test inputs are essential to systematically assess changes in accuracy across updates and thus ensure they do not lead to poorer accuracy in critical conditions.
Additionally, differential testing is helpful in detecting model reuse [12]. In industries where proprietary models are crucial assets, it is essential to verify that deployed models are original. Differential testing can identify if a model presented as new is actually derived from or modified from an existing model, thus protecting intellectual property and ensuring the authenticity of models.
In conclusion, while differential testing has broad applications, our focus in this work is specifically on model selection. Unlike other contexts, such as model compression, regression detection, or model reuse, the model selection scenario presents unique challenges. Here, the goal is not to evaluate how a single model evolves or changes, but rather to compare multiple models developed independently for the same task. These models may exhibit similar accuracy levels but differ in training data, structure, and optimization methods. Focusing on model selection allows us to explore how differential testing can expose behavioral differences between DNN models. This focus is critical because, in practice, selecting the most appropriate model, whether after training or at inference time, can directly impact the systems using these models.
@@ 2.2 Challenges
Generating triggering inputs to reveal behavioral disagreements between DNN models under test introduces three main challenges. First, uncovering discrepancies in highly accurate models is inherently difficult due to the very small input subspace leading to mispredictions. Because they are designed to perform well under a wide range of conditions, we require effective and efficient techniques to identify behavioral disagreements. Second, in many real-world scenarios, we do not have access to the internal information of DNN models. Indeed, such models are often proprietary and deployed on servers, accessible only through cloud APIs. This limitation constrains us to using black-box techniques to generate triggering inputs, as we only have access to the inputs and outputs of the DNN models. Finally, the testing budget is often limited, particularly when models (1) require extensive manual labor to label and validate triggering inputs, or (2) operate on expensive simulators or physical devices. These simulators usually provide a controlled environment that replicates true operational scenarios, allowing for rigorous testing without the costs and risks associated with testing models deployed on expensive physical devices. Within a realistic testing budget, we must effectively generate triggering inputs that reveal diverse behavioral disagreements between the models under test.
@@ 2.3 Assumptions
No Access to the Training Data: We assume that we do not have access to the original training datasets of the models under test. This is a common scenario in many real-world situations where models are proprietary or when model testing is provided as a service for example. Consequently, our ability to generate inputs and interpret model behavior is based solely on observations from models’ outputs and the currently available testing data, rather than insights from the training process. Similar Accuracy Levels: We assume that each pair of models under test yields similar and high accuracy on the currently available testing dataset. Such generation warrants the effective generation of triggering inputs to reveal behavioral disagreement between the DNN models, to better assess their respective capabilities and thus enable their use in the various practical applications mentioned above.
@@ 3.3.1 Individuals and Initial Population
In genetic search, the initial population contains individuals consisting of a set of elements called genes. These genes are connected and form an individual also referred to as a solution. In our context, the initial population contains 100 randomly generated latent vectors, each of size 100, as recommended in the literature [36]. Each dimension of a latent vector in the initial population is independently generated from a uniform distribution over the range [1,1][-1,1], thereby ensuring a diverse and unbiased initialization in the search space.
@@ 3.3.2 Fitness Functions
This fitness function estimates the divergence between the output probability vectors of the models under test, when fed with the same input image generated by the trained GAN. More specifically, the fitness function takes as input a GAN latent vector $Z$ responsible for generating the candidate test input image. We feed the models under test with such image and compute the Euclidean distance between the resulting output probability vectors $Y_{m1}$ and $Y_{m2}$. The goal of maximizing this distance is to increase the probability of selecting latent vectors that lead to the generation of input images revealing significant behavioral disagreements between the models under test.
The Euclidean distance serves as a straightforward measure of the magnitude of discrepancy between model predictions, capturing how far apart the models’ outputs are in the probability space. The choice of Euclidean distance over other metrics, such as KL divergence [37], is motivated by its efficiency in terms of computation cost.
@@
[…]
@@ 3.4.2 Mutation
Our genetic search also employs a mutation operator to further explore the search space.
Following existing works that use genetic algorithms in GAN latent space [28, 40], we also adopt polynomial mutation [42] to address the continuous nature of the latent space.
@@ 3.5 Search Algorithm
Input: Initial population of individuals ( indind $\in$ $P$ ), maximum running time ($T_{max}$), trained GAN, two DNNs under test
@@ 14    $\text{image}_{\text{gen}}\xleftarrow{}\text{GAN}(ind)$;
15    $OP1,OP2]\xleftarrow{}\text{Model1}(\text{image}_{\text{gen}}),\text{Model2}(\text{image}_{\text{gen}})$
@@ 21 $gen=gen+1$;
We set the NSGA-II parameters in our search algorithm as follows. The crossover rate CrossRateCrossRate is equal to 90% and the mutation rate MutationRateMutationRate is equal to 10%, given recommended values in the literature [44, 45, 46]. Finally, we used Google Colab and the Pymoo library [47] to implement our genetic search.